WRITE AN EQUATION OF THE LINE THROUGH THE GIVEN POINTS
(-4,-8) , (0,2)

Found 2 solutions by checkley71, Nate:
Answer by checkley71(8403)

SLOPE IS (2+8)/(0+4) OR 10/4 OR 2.5 & Y INTERCEPT IS 2 THUS
Y=2.5X+2

Answer by Nate(3500)

M = (2 + 8)/(4) = 10/4 = 5/2
y - 2 = (5/2)(x - 0)
y = (5/2)x + 2